OpenStudy (anonymous):

what is the solution to the following equation? 3 (x + 5) - 10 = -2 (x + 10) Please explain this to me!

OpenStudy (gabbyalicorn):

3(x+5)−10=−2(x+10) Simplify both sides of the equation. 3(x+5)−10=−2(x+10) Simplify: (Show steps) 3x+5=−2x−20 Add 2x to both sides. 3x+5+2x=−2x−20+2x 5x+5=−20 Subtract 5 from both sides. 5x+5−5=−20−5 5x=−25 Divide both sides by 5. 5x/5=−25/5 x=−5 Answer: x=−5

OpenStudy (anonymous):

ok thanks :)

OpenStudy (lolacole12):

Let's solve your equation step-by-step. 3(x+5)−10=−2(x+10) Step 1: Simplify both sides of the equation. 3(x+5)−10=−2(x+10) Simplify: (Show steps) 3x+5=−2x−20 Step 2: Add 2x to both sides. 3x+5+2x=−2x−20+2x 5x+5=−20 Step 3: Subtract 5 from both sides. 5x+5−5=−20−5 5x=−25 Step 4: Divide both sides by 5. 5x/5=−25/5 x=−5 Answer: x=−5

OpenStudy (lolacole12):

o wow im late

OpenStudy (anonymous):

lol

OpenStudy (lolacole12):

This should help http://www.mathpapa.com/algebra-calculator.html

OpenStudy (lolacole12):

that place is nice hehe

OpenStudy (anonymous):

lol thanks :)